OpenStudy (anonymous):

Write the equation of the circle with center (-1, 9) and containing the point (2, 5). Use the ^ key for the exponents. Write your answer like this: (x+3)^2+(y-8)^2=49

OpenStudy (anonymous):

that is easy use the equation i gave you lol.

OpenStudy (anonymous):

(x - h)2 + (y-k)2 = r2

OpenStudy (anonymous):

So that's how i would write the answer? lol

OpenStudy (anonymous):

yes plug in your knowns but you have to find the R using the distance formula.

OpenStudy (anonymous):

okay! Thanks

OpenStudy (anonymous):

remember that you are given a point on the circle and an origin and the distance is found using the distance formula.

OpenStudy (anonymous):

First, you have to know the radius of the circle. You have a point and the center, so, just use the distance formula. \[R=\sqrt{(-1-2)^2+(9-5)²} = \sqrt{9+16} = 5\] So, the equation of the circle is: \[(x-a)²+(y-k)²=R²\], where a = x coordinate of the center, k = y coordinate of the center and R = radius, so, you will have:\[(x-(-1))²+(y-(9))²=5² \iff (x+1)²+(y-9)²=25\] Answer: \[(x+1)²+(y-9)²=25\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

you're welcome :]